OpenStudy (anonymous):

help derivative find dy/dx y=5x+1/2sqx

OpenStudy (nincompoop):

this problem makes use of the quotient rule or the product rule if you know how to convert 1/x^n into n^(-n)

OpenStudy (nincompoop):

and also remember that square root in exponential form is really 1/2

OpenStudy (anonymous):

i dont got it

OpenStudy (nincompoop):

is your equation like this \[y = \frac{ 5x+1 }{ 2 \sqrt{x} }\]

OpenStudy (anonymous):

yes

OpenStudy (nincompoop):

to use the product rule with chain rule, it can be rewritten into \[y = (5x+1)[2(x^\frac{ 1 }{ 2 })]^{-1}\]

OpenStudy (nincompoop):

or if you just want to use the quotient rule, it has the form of \[y=\frac{ u }{ v }; y' = \frac{ u'v-uv' }{ v^2 }\]

OpenStudy (nincompoop):

product rule is \[y=uv \rightarrow y' = u'v+uv'\]